OpenStudy (anonymous):

(2.3 x 10^-4)(5 x 10^-11) A. 1.15 x 10^-14 B. 11.5 x 10^-15 C. 1.15 x 10^-12 D. 0.115 x 10^-14

OpenStudy (anonymous):

What is the product in scientific notation?

OpenStudy (anonymous):

2.3*5=11.5

OpenStudy (anonymous):

10^-4*10^-11=10^-15

OpenStudy (anonymous):

remember x^2*x^5=x^7 not x^10